• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Fansler/Segal (Techs)

    Log entry time 03:52:14 on November 18, 2008

    Entry number 249451

    Followups:

    keyword=Target motion controls

    Target ladder would not move up or down, rebooted the IOC, cycled poer on the velmex controller....no joy. Swapped the velmex controller, no joy. We were able to communicate directly with the velmex controller via serial connection. Replaced a bad serial cable on the Greensprings serial IP pack, no joy. We leave this problem till the day shift. The target is on the "Optics" switch, number 4. There is slop in the readback position in this case. We do not have a firm number on how much slop. It could be close to 1/4", it could be a few mm.